1.  If a random sample of 38 public elementary schools is selected , what is the probability

 That the number of students enrolled is will be more than 400 ?




1
Expert's answer
2021-06-07T16:43:18-0400

If a random sample of n=38 public elementary schools is selected, then the average number of students enrolled would be a normally distributed random variable "Y" with the mean "\mu=468" and the standard deviation "\sigma/\sqrt{n}=87/\sqrt{38}."

"P(Y>400)=1-P(Y\leq400)"

"=1-P(Z\leq\dfrac{400-468}{87/\sqrt{38}})\approx1-P(Z\leq-4.818163)"

"\approx0.9999993"